OpenStudy (anonymous):

(-1/2x).... how to find its third derivative?

OpenStudy (mathteacher1729):

Is that (1/2) times x or (1/2) to the x power? or 1 over (2x) ?

OpenStudy (anonymous):

1 over 2x

OpenStudy (anonymous):

(-1)/(2x)

OpenStudy (anonymous):

.....its 2nd derivative is 1/x^3

OpenStudy (anonymous):

@mathteacher1729

OpenStudy (anonymous):

@Callisto

OpenStudy (mathteacher1729):

Re write your original expression \[\huge f(x) = \frac{-1}{2x} =\frac{-1}{2}x^{-1}\] Use the power rule \[\huge f'(x)=(-1)\frac{-1}{2}x^{-2}=\frac{1}{2}x^{-2}\] And again... \[\huge f''(x)=(-2)\frac{1}{2}x^{-3}=-x^{-3}\] And again... :)

OpenStudy (anonymous):

@mathteacher1729 what is the power rule for that?

OpenStudy (anonymous):

numerator times the derivative of denominator? @mathteacher1729

OpenStudy (mathteacher1729):

The power rule states that \[\huge \frac{d}{dx} (x^k) = k\cdot x^{(k-1)}\]

OpenStudy (mathteacher1729):

This holds even when k is negative. :)